How do i solve this? Hydrogen and iodine gas react to form hydrogen iodide, as shown in the equation H2(g) + I2(g) → 2HI(g). Suppose you are given a time interval from t1 = 10 seconds to t2 = 20 seconds and a change in H2 concentration from0.210 mol/L to 0.185 mol/L.


1
Expert's answer
2022-06-08T14:34:04-0400

∆C = |C2–C1| = |0,185-0,21| = 0,025 mol/l

∆t = t2 – t1 = 20 - 10 = 10 sec

∆C — is the concentration change

∆t — is the time change

V — reaction rate


"V= \frac{∆C}{∆t}= \frac {0,025}{10}= 0,0025 mol/l•sec"


Answer: V= 0,0025 mol/l•sec